# DPQuant: Efficient and Differentially-Private Model Training via Dynamic Quantization Scheduling

### v1

@@ Abstract
Differentially-Private SGD (DP-SGD) is a powerful technique to protect user privacy when using sensitive data to train neural networks. During training, converting model weights and activations into low-precision formats, i.e., quantization, can drastically reduce training times, energy consumption, and cost, and is thus a widely used technique. In this work, we demonstrate that quantization causes significantly higher accuracy degradation in DP-SGD compared to regular SGD. We observe that this is caused by noise injection in DP-SGD, which amplifies quantization variance, leading to disproportionately large accuracy degradation. To address this challenge, we present DPQuant, a dynamic quantization framework that adaptively selects a changing subset of layers to quantize at each epoch. Our method combines two key ideas that effectively reduce quantization variance: (i) probabilistic sampling of the layers that rotates which layers are quantized every epoch, and (ii) loss-aware layer prioritization, which uses a differentially private loss sensitivity estimator to identify layers that can be quantized with minimal impact on model quality. This estimator consumes a negligible fraction of the overall privacy budget, preserving DP guarantees. Empirical evaluations on ResNet18, ResNet50, and DenseNet121 across a range of datasets demonstrate that DPQuant consistently outperforms static quantization baselines, achieving near Pareto-optimal accuracy-compute trade-offs and up to $2.21\times$ theoretical throughput improvements on low-precision hardware, with less than $2\%$ drop in validation accuracy.
@@ 1 Introduction
Differentially Private Stochastic Gradient Descent (DP-SGD) Abadi et al. (2016) enables training neural networks on sensitive data while providing formal privacy guarantees. To improve the efficiency of such training on modern hardware, the use of low-precision arithmetic and data formats, i.e., quantization, has gained widespread interest Gholami et al. (2021); Jacob et al. (2017). Quantization can significantly reduce the amount of computation and memory required, thus reducing the latencies, cost, and energy consumption during training and inference, often with little to no loss in model accuracy Micikevicius et al. (2022). These benefits are especially important in resource-constrained settings such as federated learning with edge devices, where support for full-precision arithmetic is limited and compute budgets are constrained.
@@
In this work, we observe that applying low-precision quantization directly to DP-SGD training often leads to significant accuracy degradation, as severe as a $40\%$ drop. While non-DP training is typically robust to quantized training, the gradient clipping and noise addition steps in DP-SGD interact poorly with low-precision arithmetic leading to poor convergence as explained in Section 4.
Our goal is to develop an automatic mechanism to effectively quantize DP-SGD, while minimizing the impact on model accuracy and the differential privacy budget. We observe that quantizing only a subset of layers and selectively varying this subset every epoch can preserve most of the efficiency gains from quantization while maintaining model accuracy. DPQuant uses two core techniques that are implemented in a differentially private framework for dynamic quantization scheduling:
@@
We make the following contributions:
@@
    To our knowledge, this work is the first to demonstrate and explain the significant accuracy degradation when employing existing quantization techniques with DP-SGD compared to non-private SGD during model training.
@@ 2 Related Works
Post-training quantization (PTQ) (Banner et al., 2019; Jacob et al., 2017; Nagel et al., 2021) aims to accelerate inference through low-precision computations. A neural network is first trained in full-precision and then its weights are quantized. The conversion to quantized formats typically involves a small calibration dataset (Hubara et al., 2021; Nagel et al., 2019) to allocate quantization bit-widths in different parts of the model and to perform bias correction. PTQ methods are orthogonal to this work since they do not optimize training.
@@
Quantization-aware training (QAT) (Krishnamoorthi, 2018) ameliorates the aforementioned accuracy loss by training in lower precision. However, this technique requires extra analysis, such as hardware simulation (Wang et al., 2019) and sensitivity estimation (Dong et al., 2019; Park et al., 2018). QAT also involves quantizer updates, such as step-size tuning (Esser et al., 2020; Ding et al., 2023) and quantizer scaling (Sakr et al., 2022), to select bit-widths (Youn et al., 2022). The incurred overhead during training typically cancels out any raw bit-width speedups and often increases wall-clock training time, making them ideal for accelerating inference but not training (Chen et al., 2024).
@@
Gradient compression (Lin et al., 2020; Alistarh et al., 2017; Alimohammadi et al., 2023; Wen et al., 2017; Shi et al., 2020) reduces communication costs by compressing gradients in distributed settings, either through sparsificiation (Stich et al., 2018; Yu et al., 2017) or low-rank approximation (Vogels et al., 2019; Idelbayev and Carreira-Perpinán, 2020). Notably, (Youn et al., 2023) combines quantization and the noising mechanism to achieve differential privacy while reducing communication. However, compression does not lower the arithmetic cost of training. In addition, these methods often rely on assumptions such as full gradient availability or error feedback accumulation (Karimireddy et al., 2019), which are difficult to satisfy under DP constraints.
Mixed-precision training Choi et al. (2018); Zhou et al. (2018); Sun et al. (2020); Chmiel et al. (2024); Micikevicius et al. (2022) aims to reduce training cost by operating on lower-precision data types, e.g., FP16, BF16, FP8, and FP4). While effective for standard SGD, mixed-precision training degrades significantly under DP-SGD. To our knowledge, no prior work has explored mixed-precision training when differential privacy mechanisms are employed.
@@ 3.1 Differentially-Private DNN Training
Definition 1 (Differential Privacy, Dwork and Roth (2014)).
@@
$\theta_{t+1}\leftarrow\theta_{t}-\eta\left(\frac{1}{|B|}\sum_{i\in B}\mathrm{clip}(\nabla\mathcal{L}(\theta_{t},x_{i}))+\mathcal{N}(0,\sigma^{2}C^{2}\mathds{1})\right),$
where $B$ is a minibatch of training examples, $\mathrm{clip}(\cdot)$ scales the gradient to have $\ell_{2}$ norm at most $C$, and $\mathcal{N}(0,\sigma^{2}C^{2}\mathds{1})$ is Gaussian noise added to ensure privacy.
@@ 3.2 Quantization and Mixed Precision Training
Modern hardware accelerators such as NVIDIA GPUs, Google TPUs, and Qualcomm Hexagon NPUs provide dedicated support for low-precision arithmetic, including fp16, bfloat16, and increasingly lower bitwidth formats like fp8, fp6, and fp4. These formats enable faster matrix multiplications and convolutions by reducing arithmetic complexity, memory usage, and data transfer costs. Lower precision reduces both the number of transistors required per operation and the bandwidth needed for memory and interconnects, resulting in substantial speedups and energy savings.
While prior work has demonstrated that full training in low-bit formats (e.g., fp4) can retain accuracy under standard SGD, extending these techniques to differentially private training remains challenging. The clipping and noise injection steps in DP-SGD amplify quantization errors and increase gradient variance, making DP training more sensitive to precision loss. Fully quantized DP-SGD thus often results in severe degradation unless carefully tuned.
4 Degradation of Quantized DP-SGD
@@
Figure 1: Comparing quantized SGD vs DP-SGD ResNet18 training on the GTSRB dataset
Figure 1 presents a case study of ResNet18 trained on GTSRB, where the forward and backward convolution operators are quantized to evaluate the effects of quantized training. Figure 1(a) shows the accuracy loss compared to the unquantized baseline for different degrees of quantization (in terms of the number of layers quantized), and the error bars represent the results when different subsets of layers are chosen for quantization, both for DP and non-DP training. For the non-DP SGD baseline, fully-quantized training results in only a modest accuracy drop of around 1%. In contrast, DP-SGD experiences a much greater degradation, up to 5%. Furthermore, the variance in performance due to different layers being quantized is substantially higher under DP-SGD. We observe similar trends in other neural networks and datasets (included in Appendix A.2).
We hypothesize that the increased sensitivity to quantization can be attributed to noise injection in DP-SGD as follows. In iteration $t$ in the DP-SGD training, the gradients $\mathbf{g}_{t}$ is first clipped to obtain $\bar{\mathbf{g}}_{t}$ where $\bar{\mathbf{g}}_{t}||_{2}\leq C$ (section 3.1). Next, noise $\mathbf{n}_{t}\sim\mathcal{N}(0,\sigma^{2}C^{2}\mathds{1})$ is sampled, and finally the weights are updated using the sum of the noise and clipped gradient:
@@
$\|\mathbf{n}_{t}\|_{\infty}\;\approx\;\|\bar{\mathbf{g}}_{t}\|_{2}\;\gg\;\|\bar{\mathbf{g}}_{t}\|_{\infty}$ (2)
@@
where the asymptotic equality holds due to Equation 2. Assuming $L$-Lipschitz-smoothness of the loss with respect to the gradients:
@@
We now show that the $\infty$-norm of the raw gradients (i.e. before clipping and noising) of the next iteration is bounded by $\||\bar{\mathbf{g}}_{t}||_{2}$ using the inverse triangle inequality with Equation 4 and 5:
@@
To show this empirically, we plot the norms of intermediate gradients under both SGD and DP-SGD using the same hyperparameters in Figure 1(c), where the DP-SGD intermediate raw gradients are $2\times$ larger in the average and worst case. This phenomenon has also been observed in prior work Du et al. (2022), showing a even larger gap than what we observe in gradient norms later in training.
@@
(under DP-SGD) $\qquad\mathrm{Var}\bigl{(}q(\mathbf{g}_{t+1})\,|\,\mathbf{g}_{t+1}\bigr{)}=\mathcal{O}\bigl{(}\|\mathbf{g}_{t+1}\|_{\infty}^{2}\bigr{)}=\mathcal{O}\bigl{(}\|\mathbf{g}_{t}\|_{2}^{2}\bigr{)}$
(under SGD) $\qquad\mathrm{Var}\bigl{(}q(\mathbf{g}_{t+1})\,|\,\mathbf{g}_{t+1}\bigr{)}=\mathcal{O}\bigl{(}\|\mathbf{g}_{t}\|_{\infty}^{2}\bigr{)}$
The quantization variance above is in addition to the existing variance of the stochastic gradients, as well as noise injected by DP-SGD. In higher dimensions, $\|\mathbf{g}_{t}\|_{2}\gg\|\mathbf{g}_{t}\|_{\infty}$, quantization contributes much more variance to the gradients, hence leads to slower and less reliable convergence Johnson and Zhang (2013) and accuracy degradation. To address this challenge, we aim to reduce the quantized-induced variance.
@@ 5.1 Part I: Probabilistic Layer Sampling
Suppose a layer is quantized with probability $p$, we let $\mathbf{g}_{\text{fp}}$ to denote its full precision gradients and $\mathbf{g}_{\text{quant}}$ to be its gradients computed under quantization. By Section 4, quantization incurs additional variance, hence $\operatorname{Var}(\mathbf{g}_{\text{fp}})\leq\operatorname{Var}(\textbf{g}_{\text{quant}})$. We can write the expected gradient variance as:
From this it follows that whenever $p<1$ – that is, when only a subset of layers is quantized at each epoch—the average quantization-induced variance is strictly lower than in full quantization. Furthermore, by rotating which layers are quantized every epoch, no single layer repeatedly incurs the full quantization variance, and hence their expected variance remains smaller than $\operatorname{Var}(\mathbf{g}_{\text{quant}})$.
@@ 5.2 Part II: Loss-Aware Layer Prioritization
Our goal is to find policies that minimally increases loss (or equivalently, a policy $p$ with small $R(p)$). We note that a key challenge in evaluating $R(p)$ for a given policy $p$ is that the expectation is taken over the full private dataset $D$. This makes direct computation both expensive and incompatible with tight privacy guarantees. To address this, we instead estimate $R(p)$ by subsampling $D$ and running a limited number of DP-SGD iterations under policy $p$ to obtain a proxy loss, then the same training iterations is done to obtain the baseline full-precision loss, and the difference is used as an empirical estimate of the loss impact.
@@
5.3 Dynamic Layer Selection for Quantized DP-SGD Training
Building on the insights from the previous sections, we design a dynamic layer selection strategy for quantized DP-SGD that combines: (i) probabilistic sampling of quantized layers to reduce variance (Section 5.1), and (ii) loss-aware prioritization to preserve performance by avoiding quantization of high-impact layers (Section 5.2).
Table 1: DPQuant hyperparameters
Parameter Default Description
$n$ 60 number of epochs to train
$k$ layers to execute in low precision.
$n_{\text{sample}}$ 1 test samples for loss measurement.
$n_{\text{interval}}$ 2 epochs to train before the next measurement.
$R$ 2 repetitions during measurement.
$\sigma_{\text{measure}}$ 0.5 Noise scale used during loss-difference privatization.
$\mathcal{C}_{\text{measure}}$ 0.01 Clipping norm used during loss-difference privatization.

### v2

@@ Abstract
Differentially-Private SGD (DP-SGD) and its adaptive variant DP-Adam are powerful techniques to protect user privacy when using sensitive data to train neural networks. During training, converting model weights and activations into low-precision formats, i.e., quantization, can drastically reduce training times, energy consumption, and cost, and is thus a widely used technique. In this work, we demonstrate for the first time that quantization causes significantly higher accuracy degradation in DP training compared to regular SGD. We observe that this is caused by noise injection, which amplifies quantization variance, leading to disproportionately large accuracy degradation. To address this challenge, we present DPQuant, a dynamic quantization framework that adaptively selects a changing subset of layers to quantize at each epoch. Our method combines two key ideas that effectively reduce quantization variance: (i) probabilistic sampling that rotates which layers are quantized every epoch, and (ii) loss-aware layer prioritization, which uses a differentially private loss sensitivity estimator to identify layers that can be quantized with minimal impact on model quality. This estimator consumes a negligible fraction of the overall privacy budget, preserving DP guarantees. Empirical evaluations on ResNet18, ResNet50, and DenseNet121 across a range of datasets demonstrate that DPQuant consistently outperforms static quantization baselines, achieving near Pareto-optimal accuracy-compute trade-offs and up to $2.21\times$ theoretical throughput improvements on low‑precision hardware, with less than $2\%$ drop in validation accuracy. We further show that our framework extends to DP-Adam with similar gains.
@@ 1 Introduction
Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al., 2016) enables training neural networks on sensitive data while providing formal privacy guarantees. To improve the efficiency of such training on modern hardware, the use of low-precision arithmetic and data formats, i.e., quantization, has gained widespread interest (Gholami et al., 2021, Jacob et al., 2017). Quantization can significantly reduce the amount of computation and memory required, thus reducing the latencies, cost, and energy consumption during training and inference, often with little to no loss in model accuracy (Micikevicius et al., 2022). These benefits are especially important in resource-constrained settings such as federated learning with edge devices, where support for full-precision arithmetic is limited and compute budgets are constrained.
@@
In this work, we observe that applying low-precision quantization directly to DP training often leads to significant accuracy degradation, as severe as a $40\%$ drop. While non-DP training is typically robust to quantization, the gradient clipping and noise addition steps in DP-SGD/DP-Adam interact poorly with low-precision arithmetic leading to poor convergence as explained in Section 4.
Our goal is to develop an automatic mechanism to effectively quantize DP training, while minimizing the impact on model accuracy and the privacy budget. We observe that quantizing only a subset of layers and selectively varying this subset every epoch can preserve most of the efficiency gains from quantization while maintaining model accuracy. DPQuant uses two core techniques that are implemented in a differentially private framework for dynamic quantization scheduling:
@@
Specifically, in this work we make the following contributions:
@@
    We are the first to demonstrate and explain the significant accuracy degradation that arises when employing ultra-low-precision (FP4/INT4) quantization with DP training (DP-SGD and DP-Adam), compared to the mild impact observed in non-private SGD and in 16-bit DP training.
@@ 2 Related Works
Post-training quantization (PTQ) (Banner et al., 2019, Jacob et al., 2017, Nagel et al., 2021) aims to accelerate inference through low-precision computations. A neural network is first trained in full-precision and then its weights are quantized. The conversion to quantized formats typically involves a small calibration dataset (Hubara et al., 2021, Nagel et al., 2019) to allocate quantization bit-widths in different parts of the model and to perform bias correction. PTQ methods are orthogonal to this work since they do not optimize training.
@@
Quantization-aware training (QAT) (Krishnamoorthi, 2018) ameliorates the aforementioned accuracy loss by training in lower precision. However, QAT requires additional overhead such as hardware simulation (Wang et al., 2019), sensitivity estimation (Dong et al., 2019), and quantizer tuning (Esser et al., 2020, Sakr et al., 2022), which typically cancels out raw bit-width speedups, making QAT suited for accelerating inference but not training (Chen et al., 2024).
@@
Gradient compression (Lin et al., 2020, Alistarh et al., 2017, Alimohammadi et al., 2023, Wen et al., 2017, Shi et al., 2020) reduces communication costs by compressing gradients in distributed settings, either through sparsification (Stich et al., 2018, Yu et al., 2017) or low-rank approximation (Vogels et al., 2019, Idelbayev and Carreira-Perpinán, 2020). Notably, Youn et al. (2023) combine quantization and the noising mechanism to achieve differential privacy while reducing communication. However, compression does not lower the arithmetic cost of training. In addition, these methods often rely on assumptions such as full gradient availability or error feedback accumulation (Karimireddy et al., 2019), which are difficult to satisfy under DP constraints.
Mixed-precision training (Choi et al., 2018, Zhou et al., 2018, Sun et al., 2020, Chmiel et al., 2024, Micikevicius et al., 2022) aims to reduce training cost by operating on lower-precision data types (e.g., FP16, BF16, FP8, and FP4). Prior work has shown that 16-bit mixed-precision training (FP16/BF16) is compatible with DP-SGD with little accuracy loss (Li et al., 2022, Bu et al., 2023). However, pushing to ultra-low precision (FP4/INT4) under DP remains challenging: we show in Section 4 that DP noise amplifies quantization variance at these bit-widths, causing significant accuracy degradation that does not arise at 16-bit precision.
Quantization as a privacy mechanism. Several works use quantization to provide or strengthen DP guarantees in federated or decentralized settings (Kang et al., 2024, Kim et al., 2022, Wang and Basar, 2022, Xiong et al., 2016). In these approaches, quantization is applied after gradients are computed in full precision, so there is no effect on subsequent forward/backward passes. DPQuant instead quantizes activations and weights during training to reduce arithmetic cost.
Improved DP mechanisms. A complementary line of work improves the DP algorithm itself for better utility at a given $\varepsilon$, whether through modified optimizers (Wang et al., 2024, Wei et al., 2022, Park et al., 2023, Denisov et al., 2023) or by reducing the trainable parameter space via pruning or layer selection (Adamczewski and Park, 2023). These methods are orthogonal to DPQuant, which keeps the full dense model and instead optimizes how much computation can run in low precision.
Gradient clipping optimizations for DP-SGD. (Li et al., 2022, Bu et al., 2022, Lee and Kifer, 2020, Subramani et al., 2021) optimize the per-sample gradient clipping for DP-SGD by eliminating redundant computation and increasing vectorized computation, leading to better hardware utilization. These works are orthogonal to DPQuant.
@@ 3.1 Differentially-Private DNN Training
Definition 1 (Differential Privacy, (Dwork and Roth, 2014)).
@@
$\theta_{t+1}\leftarrow\theta_{t}-\eta\left(\frac{1}{|B|}\sum_{i\in B}\mathrm{clip}_{C}(\nabla\mathcal{L}(\theta_{t},x_{i}))+\mathcal{N}(0,\sigma^{2}C^{2}\mathds{1})\right),$
where $B$ is a minibatch of training examples, $\mathrm{clip}_{C}(\cdot)$ scales the gradient to have $\ell_{2}$ norm at most $C$, and $\mathcal{N}(0,\sigma^{2}C^{2}\mathds{1})$ is Gaussian noise added to ensure privacy.
@@ 3.2 Quantization and Mixed Precision Training
Modern hardware accelerators such as NVIDIA GPUs, Google TPUs, and Qualcomm Hexagon NPUs provide dedicated support for low-precision arithmetic, including fp16, bfloat16, and increasingly lower bitwidth formats like fp8, fp6, and fp4. These formats enable faster computation by reducing arithmetic complexity, memory footprint, and data transfer costs.
While prior work(Chmiel et al., 2024) has demonstrated that full training in low-bit formats (e.g., fp4) can often retain accuracy under standard SGD, extending these techniques to differentially private training remains challenging. The clipping and noise injection steps in DP-SGD amplify quantization errors and increase gradient variance, making DP training more sensitive to precision loss. Fully quantized DP-SGD thus often results in severe degradation unless carefully tuned.
4 Degradation of DP Training from Quantization
@@
Figure 1: Comparing quantized SGD vs DP-SGD ResNet18 training on the German Traffic Sign Recognition Benchmark (GTSRB) (Stallkamp et al., 2011) dataset
Figure 1 presents a case study of ResNet18 trained on GTSRB, where the forward and backward convolution operators are quantized to FP4 to evaluate the effects on training. Figure 1(a) shows the accuracy loss compared to the non-quantized baseline for different degrees of quantization (in terms of the number of layers quantized), and the error bars represent the results when different subsets of layers are chosen for quantization, both for DP and non-DP training. For the non-DP SGD baseline, fully-quantized training results in only a modest accuracy drop of around 1%. In contrast, DP-SGD experiences a much greater degradation, up to 5%. Furthermore, the variance in performance due to different layers being quantized is substantially higher under DP-SGD. We observe similar trends in other neural networks and datasets (included in Appendix A.6).
We hypothesize that the increased sensitivity to quantization can be attributed to noise injection in DP-SGD as follows. In iteration $t$ in the DP-SGD training, the gradients $\mathbf{g}_{t}$ are first clipped to obtain $\bar{\mathbf{g}}_{t}$ where $\bar{\mathbf{g}}_{t}||_{2}\leq C$ (section 3.1). Next, noise $\mathbf{n}_{t}\sim\mathcal{N}(0,\sigma^{2}C^{2}\mathds{1})$ is sampled, and finally the weights are updated using the sum of the noise and clipped gradient:
@@
[…]
@@
where the asymptotic equality holds due to Equation 2. Assuming $L$-Lipschitz‐smoothness of the loss with respect to the gradients:
@@
We now show that the $\infty$-norm of the raw gradients (i.e. before clipping and noising) of the next iteration is bounded by $\bar{\mathbf{g}}_{t}||_{2}$ using the inverse triangle inequality with Equation 4 and 5:
@@
Notably, the batch size has a negligible effect on norms, even though a larger batch size leads to a smaller variance of the stochastic gradients. Therefore, we omit it in this analysis; we include further discussions and empirical evaluations of this in appendix A.1.
To show this empirically, we plot the norms of intermediate gradients under both SGD and DP-SGD using the same hyperparameters in Figure 1(c), where the DP-SGD intermediate raw gradients are $2\times$ larger in the average and worst case. This phenomenon has also been observed in prior work (Du et al., 2022), showing an even larger gap than what we observe in gradient norms later in training.
@@
[…]
Extension to DP-Adam/DP-AdamW.
The above argument for DP-SGD relies on DP noise dominating the raw gradients, which leads to disproportionate quantization variance. In DP-Adam, noise is added to the clipped gradients first, and Adam’s preconditioner then rescales the noisy gradients coordinate-wise. Since the preconditioner applies the same scaling to both signal and noise, it preserves the relative scale between them, so quantized DP-Adam incurs the same quantization overhead as DP-SGD.
We can reason about this more directly by considering the per-coordinate signal-to-noise ratio. In DP-Adam, each noisy gradient coordinate is $\tilde{g}_{i}=g_{i}+n_{i}$, where $g_{i}$ is the true gradient and $n_{i}\sim\mathcal{N}(0,\sigma^{2}C^{2})$. The preconditioner divides by a running second-moment estimate, so at stationarity the preconditioned coordinate is $u_{i}\approx(g_{i}+n_{i})/\sqrt{g_{i}^{2}+\sigma^{2}C^{2}}$ with $\operatorname{Var}(u_{i})\leq 1$.
[…]
$|\mathbb{E}[u_{i}]|=\frac{|g_{i}|}{\sqrt{g_{i}^{2}+\sigma^{2}C^{2}}}\approx\frac{|g_{i}|}{\sigma C}\leq\frac{1}{\sigma}$ (7)
[…]
$\text{SNR}_{i}\coloneqq\frac{(\mathbb{E}[u_{i}])^{2}}{\operatorname{Var}(u_{i})}\approx\frac{g_{i}^{2}/(\sigma^{2}C^{2})}{1}=\frac{g_{i}^{2}}{\sigma^{2}C^{2}}.$ (8)
Thus the SNR in DP-Adam matches that of DP-SGD. Therefore, DP turns each preconditioned coordinate into unit-variance noise with very small mean. A fixed-precision quantizer on $u_{i}$ then adds zero-mean noise whose variance does not depend on this small mean, so the signal-to-quantization-noise ratio behaves like $g_{i}^{2}/(\sigma^{2}C^{2})$ and deteriorates as privacy noise increases.
@@ 5.1 Part I: Probabilistic Layer Sampling
Suppose a layer is quantized with probability $p$, we let $\mathbf{g}_{\text{fp}}$ denote its full precision gradients and $\mathbf{g}_{\text{quant}}$ to be its gradients computed under quantization. By Section 4, quantization incurs additional variance, hence $\operatorname{Var}(\mathbf{g}_{\text{fp}})\leq\operatorname{Var}(\mathbf{g}_{\text{quant}})$. The expected gradient variance is:
@@
From this it follows that whenever $p<1$ – that is, when only a subset of layers is quantized at each epoch – the average quantization‐induced variance is strictly lower than in full quantization. Furthermore, by rotating which layers are quantized every epoch, no single layer repeatedly incurs the full quantization variance, and hence their expected variance remains smaller than $\operatorname{Var}(\mathbf{g}_{\text{quant}})$.
@@ 5.2 Part II: Loss-Aware Layer Prioritization
Our goal is to find policies that minimally increase loss (or equivalently, a policy $p$ with small $R(p)$). We note that a key challenge in evaluating $R(p)$ for a given policy $p$ is that the expectation is taken over the full private dataset $D$. This makes direct computation both expensive and incompatible with tight privacy guarantees. To address this, we instead estimate $R(p)$ by subsampling $D$ and running a limited number of DP-SGD iterations under policy $p$ to obtain a proxy loss, then the same training iterations are done to obtain the baseline full-precision loss, and the difference is used as an empirical estimate of the loss impact.
@@
5.3 Dynamic Layer Selection for Quantized DP Training
Building on the insights from the previous sections, we design a dynamic layer selection strategy for quantized DP training that combines: (i) probabilistic sampling of quantized layers to reduce variance (Section 5.1), and (ii) loss-aware prioritization to preserve performance by avoiding quantization of high-impact layers (Section 5.2).